OpenStudy (anonymous):

Eliminate the “z” variable in the first two equations by finding opposite coefficients and adding. What is the resulting equation? 4x – y + z = 6 3x + 2y – z = 0 4x – y – 3z = 2

OpenStudy (anonymous):

Whenever i'm solving for a linear system of equations, i usually number my equations to keep track of them just in case. Here, lets order them from top to bottom, 1, 2 and 3. So if we add, 1+2 we get 4x-y+z = 6 + 3x+2y-z=0 -------------- Now just add like terms. basically add straight down. What do you get?

OpenStudy (anonymous):

thnx